OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is -11, the last term is -45,056, and the common ratio is -4? -143,231 -36,047 144,177 716,144

OpenStudy (solomonzelman):

I had to refersh sorry

OpenStudy (solomonzelman):

\(\large\color{black}{ S_n=\frac{\LARGE a_1(1-r^n)}{\LARGE 1-r}}\)

OpenStudy (anonymous):

What would n be?

OpenStudy (solomonzelman):

you know that: ~ \(\large\color{black}{ a_1 =-11}\) ~ \(\large\color{black}{ r =-4}\) ~ \(\large\color{black}{ n =7}\) ~ \(\large\color{black}{ S_n =S_7}\)

OpenStudy (anonymous):

-36,047

OpenStudy (solomonzelman):

that is the sum?

OpenStudy (anonymous):

Wouldn't it be B because if you fill out the equation and find the answer I got -36.047

OpenStudy (solomonzelman):

I'll check this

OpenStudy (solomonzelman):

\(\large\color{black}{ S_n=\frac{\LARGE (-11)(1-(-4)^7)}{\LARGE 1-7}}\) \(\large\color{black}{ S_n=\frac{\LARGE (-11)(1+4^7)}{\LARGE 6}}\)

OpenStudy (solomonzelman):

6 is negtative

OpenStudy (solomonzelman):

(-11)(1+4^7)/-6

OpenStudy (solomonzelman):

(11)(1+4^7)/6

OpenStudy (jhannybean):

Hitn: they give you the last term: -45,056

OpenStudy (anonymous):

r=-4 so if you put 1-r it would turn into 1+4 meaning divide by +5 right?

OpenStudy (solomonzelman):

you don't need the 7th term for the sum

OpenStudy (solomonzelman):